Please help me sokve this equation; 5(3+5x)/6=3/7 what is x?
Answer by algebrahouse.com(1659)

5(3 + 5x)/6 = 3/7
5(3 + 5x) = 18/7 {multiplied both sides by 6}
15 + 25x = 18/7 {used distributive property}
25x = -87/7 {subtracted 15 from both sides}
x = -87/175 {divided both sides by 25, multiplying by 1/25}

For more help from me, visit: www.algebrahouse.com